OpenStudy (anonymous):

tan3x-tan x= tan2x ....? is it true?

OpenStudy (callisto):

since tan(A-B) = (tanA -tanB)/ ( 1+ tanAtanB) (tanA -tanB) = tan (A-B) ( 1+ tanAtanB ) tan 3x -tanx = tan (3x-x) (1+ tan3xtanx) = tan 2x ( 1+tan3xtanx) So... what do you think?